Question 20 In the 2012 presidential election, exit polls from the critical state of Ohio provided the following results: Total Obama Romney 4590 no college degree (60%) college graduate (40%) 52% 47% 51% What is the probability a randomly selected respondent voted for Obama? Give the exact answer. exact number, no tolerance
Answer #1

Given that,

P(Obama|No college degree) = 52/100 * 60/100= 0.312 & P(Romney|No colllege degree) = 45/100 * 60/100 = 0.270

P(Obama|College degree) = 47/100 * 40/100= 0.188 & P(Romney|College degree) = 51/100 * 40/100= 0.204

If a voter select randomly. then if he may from college degree or No college degree.

So, P(Obama)= P(Obama|No college degree) + P(Obama|College degree) = 0.312+0.188= 0.5